Express each of the following as single trigonometric ratio

Use double angle formula

a) 2 sin 5x cos 5x
b) cos^2 0 - sin ^2 0
c) 1- 2 sin^2 3x
d) 2 tan 4x / 1-tan^2 4x
1
Expert's answer
2020-10-27T19:47:11-0400

a) 2 sin 5x cos 5x=sin10x

b) cos2 0 - sin2 0=cos(2*0)

c) 1- 2 sin2 3x=cos6x

d) 2 tan 4x / (1-tan2 4x)=tan8x